OpenStudy (anonymous):

2. The total number of degrees in a regular hexagon is 720°. If a corresponding angle in a congruent regular hexagon is (11x – 12)°, find the value of x. 2. The total number of degrees in a regular hexagon is 720°. If a corresponding angle in a congruent regular hexagon is (11x – 12)°, find the value of x. @Mathematics

hero (hero):

720/6 = 120 120 = 11x-12 120 + 12 = 11x 132 = 11x 132/11 = x 12 = x

hero (hero):

Sorry, I get too excited sometimes and make mistakes

OpenStudy (anonymous):

Hero check out what I responded

OpenStudy (anonymous):

Sorry to interrupt